OpenStudy (anonymous):

What is the product of nine over ten times two and two over three in simplest form? two and eighteen over thirty two and three over ten two and eleven over thirteen two and two over five

OpenStudy (anonymous):

what kind of question is that?

OpenStudy (anonymous):

\[\frac{9}{10} \times 2 \times \frac{2}{3}=\] you have to multiply and reduce.

OpenStudy (anonymous):

I TRULY SUCK AT FRACTIONS

OpenStudy (anonymous):

wait... how?

OpenStudy (anonymous):

9*2=18 10*2=20

OpenStudy (anonymous):

18/20 x 2/3

OpenStudy (anonymous):

18*2=36 20*3=60

OpenStudy (anonymous):

36/60 / 1/2

OpenStudy (anonymous):

\[\frac{9}{10} \times 2 \times \frac{2}{3}= \frac{9}{10} \times \frac{2}{1} \times \frac{2}{3}= \frac{9 \times 2 \times 2}{10 \times 1 \times 3}=?\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

what is right, your answer? no it is not.

OpenStudy (anonymous):

36/60? or 36/30?

OpenStudy (anonymous):

Look at my work above.

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

Do you see what I did before? I am getting \[\frac{9 \times 2 \times 2}{10 \times 1 \times 3}=?\]

OpenStudy (anonymous):

I did and I got this: 10*1*3=30 9*2*2

OpenStudy (anonymous):

it is a

OpenStudy (anonymous):

No it is not A.

OpenStudy (anonymous):

\[\frac{36}{30}= \frac{36 \div 6 }{30 \div 6}=?\]

OpenStudy (anonymous):

see how it is 36/30?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

DUDE!!! it is A

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

I HAS TO GO PLEASE TELL ME THE ANSWER!

OpenStudy (anonymous):

I'VE BEEN HERE FOR A LONG TIME!

OpenStudy (anonymous):

I misread the problem, sorrry!

OpenStudy (anonymous):

is it a or not???

OpenStudy (anonymous):

PLEASE answer

OpenStudy (anonymous):

\[\frac{9}{10}\times 2\frac{2}{3}=\frac{9}{10}\times \frac{2\times3 + 2}{2}=\frac{9}{10}\times\frac{8}{2}=\frac{9\times8}{10 \times 2}=\frac{72}{10}=\frac{36}{5}\]

OpenStudy (anonymous):

you can simplify that!

OpenStudy (anonymous):

simplify what? the fraction? not simplify it, but write it as a mixed number. \[7\frac{1}{5}\]

OpenStudy (anonymous):

But it is not any f the answers, no matter how you read the problem

OpenStudy (anonymous):

yeah yeah... cool and everything... but that is not an answer

OpenStudy (anonymous):

@Objection, @Directrix can you show me my mistake?

OpenStudy (anonymous):

I personally think it is a

OpenStudy (anonymous):

Hold on, don't go, the question, might be wrong, or it might be just any other chiose.

OpenStudy (anonymous):

okay... I will wait

OpenStudy (anonymous):

i will call my friends also...

OpenStudy (anonymous):

@hartnn help! @hihi67 plz come guys... ur the best!

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

he left

OpenStudy (anonymous):

@hihi67 PLEASE COME!!!!!

OpenStudy (hihi67):

What do you need?

OpenStudy (anonymous):

OK UR HERE!

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

don't u think it is a?

OpenStudy (hihi67):

Yes

OpenStudy (anonymous):

okay! finnaly!

OpenStudy (anonymous):

thank you bro!

OpenStudy (hihi67):

...what?

OpenStudy (anonymous):

@hihi67, can you wirt ei tout, I am confused.

OpenStudy (anonymous):

write it out...

OpenStudy (hihi67):

WHY?

OpenStudy (anonymous):

I want to see.

OpenStudy (anonymous):

cause he doasnt get it!

OpenStudy (anonymous):

were all confused!

OpenStudy (hihi67):

Yeah...no thanks.

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

How would you know the answer, if you can't write it out?

OpenStudy (hihi67):

I know the answer, it takes forever to type it though.

OpenStudy (anonymous):

@hartnn come help us!

OpenStudy (anonymous):

okay so ppl... IS IT A OR NOT?!?!??!?

OpenStudy (hihi67):

It's a, no get on with your life.

OpenStudy (anonymous):

OKAY! THANK YOU!

OpenStudy (anonymous):

BYE PPL

OpenStudy (hihi67):

np

OpenStudy (anonymous):

How do you know the answer is a?

OpenStudy (anonymous):

just get over it! okay i will write it

OpenStudy (hihi67):

ikr

OpenStudy (anonymous):

Lets start over, @hartnn, can you write this out for us? What is the product of nine over ten times two and two over three in simplest form? two and eighteen over thirty two and three over ten two and eleven over thirteen two and two over five

OpenStudy (anonymous):

i got to the conclusion 36/60

OpenStudy (anonymous):

and when u simplify it... u get a!

OpenStudy (anonymous):

@carolinagb02, where do you get 36/60 from?

hartnn (hartnn):

i was going over the comments....what Help_me have written is correct

OpenStudy (anonymous):

Yeah, both ways it is not any of the choices no matter how you write it out.

OpenStudy (anonymous):

okat so u remember u did 10*1*3?

OpenStudy (anonymous):

that was wrong!

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

@carolinagb02, how would you write the problem out, before saying I am right or wrong?

OpenStudy (anonymous):

get 9 and multiply by 2

OpenStudy (anonymous):

u get 18

OpenStudy (anonymous):

Hold on, can you writ the problem out, the way you think?

hartnn (hartnn):

it could be this too 9/ (10*2) * 2/3 = 3/10 hmm

OpenStudy (anonymous):

10*3 and 9*2

OpenStudy (anonymous):